OpenStudy (anonymous):

Little Susie is enjoying a nice spherical lollipop. She sucks the lollipop in such a way that the circumference decreases by 1 centimeter per minute. How fast is the volume of her lollipop changing when the lollipop has a radius of 5 centimeters?

OpenStudy (anonymous):

V = (4/3)pi r^3 dV/dt = 4pi r^2 (dr/dt) dC/dt = 2pi (dr/dt) dr/dt = (1/2pi) dC/dt dV/dt = (4pi/2pi)(5)^2 (-1) dV/dt = -2*5^2*1 = - 50 cubic centimeters/minute

OpenStudy (anonymous):

ho-leee crap. a ten centimeter wide spherical lollipop? i gotta get me one of those... :P